### v2

# CLIPLoss and Norm-Based Data Selection Methods for Multimodal Contrastive Learning

Abstract
Data selection has emerged as a core issue for large-scale visual-language model pretaining (e.g., CLIP), particularly with noisy web-curated datasets. Three main data selection approaches are: (1) leveraging external non-CLIP models to aid data selection, (2) training new CLIP-style embedding models that are more effective at selecting high-quality data than the original OpenAI CLIP model, and (3) designing better metrics or strategies universally applicable to any CLIP embedding without requiring specific model properties (e.g., CLIPScore is one popular metric). While the first two approaches have been extensively studied, the third remains under-explored. In this paper, we advance the third approach by proposing two new methods. Firstly, instead of classical CLIP scores that only consider the alignment between two modalities from a single sample, we introduce negCLIPLoss, a CLIP loss-inspired method that adds the alignment between one sample and its contrastive pairs as an extra normalization term for better quality measurement. Secondly, when downstream tasks are known, we propose a new norm-based metric, NormSim, to measure the similarity between pretraining data and target data. We test our methods on the data selection benchmark, DataComp (gadre2023datacomp). Compared to the best baseline using only OpenAI’s CLIP-L/14, our methods achieve a 5.3% improvement on ImageNet-1k and a 2.8% improvement on 38 downstream evaluation tasks. Moreover, both negCLIPLoss and NormSim are compatible with existing techniques. By combining our methods with the current best methods DFN (fang2023data) and HYPE (kim2024hype), we can boost average performance on downstream tasks by 0.9%, achieving a new state-of-the-art.
Keywords: contrastive learning, visual-language pretraining, data selection, CLIP
1 Introduction
Curating large-scale visual-language datasets from web-sourced data has become common for pretraining multi-modal models. However, the quality of these web-curated data pairs remains a critical bottleneck. Research has shown that the choice of dataset significantly impacts model performance, irrespective of the models and training techniques employed (radford2021learning; schuhmann2022laion; cherti2023reproducible; bai2023qwen), and this motivates the development of various data selection strategies. This paper focuses on optimizing subset selection from a fixed data pool to train a CLIP model (radford2021learning) that achieves superior performance on zero-shot downstream tasks.
Classical methods rely solely on OpenAI’s pretrained CLIP model (i.e., a teacher model) and focus on better utilizing the embeddings. The most commonly used one is calculating CLIPScore (hessel2021clipscore), which measures the cosine similarity between the visual and language embeddings of the CLIP model for the same sample, to eliminate low-quality data with mismatches between text and image. Other works also leverage heuristic distribution alignment techniques to select samples relevant to downstream tasks, such as image-based filtering (gadre2023datacomp). These approaches are generally viewed as providing only limited enhancements. However, we argue that the potential of those embeddings has been heavily under-explored. This work seeks a universal method to better employ any given embeddings, not only from OpenAI’s (OAI) CLIP, but also from other CLIP-style models.
On the other hand, recent leading data filtering methods, instead of focusing on improving embedding utilization stategy itself, mainly follow the other two directions, both employing external resources. They either (1) use external non-CLIP models that aid in data selection, (2) or use external high-quality multi-modal data to train a better CLIP-style embedding model than the original OAI CLIP to filter out low-quality data. Specifically, in the first line of works, HYPE (kim2024hype) leverages embeddings from hyperbolic models instead of the classical Euclidean-based CLIP to measure how each data point has semantically overlaps with other data points and filters out data with low specificity. T-MARS (maini2023t) removes images where the text is the only feature correlated with the caption using FAST (chen2021fast), an off-the-shelf OCR text detection model. Devil (yu2023devil) applies fasttext (joulin2016bag) to remove non-English texts and use BLIP-2 (li2023blip) model for digit recognition to keep useful images with digits. The second direction, represented by Data Filtering Network (DFN) (fang2023data), involves training a new CLIP-style teacher model that uses high-quality datasets like HQITP-350M. Although the embeddings extracted from this model perform worse than the OAI CLIP in downstream tasks, it is particularly good at filtering out low-quality data. Notably, some of these methods can be combined and indeed, merging the selected data from DFN and HYPE achieves current state-of-art as shown in HYPE paper (kim2024hype).
In this work, we focus on improving the filtering methods themselves for any given CLIP embedding rather than mainly relying on external models or multi-modal datasets. We show that there are universal and more effective strategies for utilizing any CLIP teacher model, regardless of its architecture (e.g., B/32 or L/14) or the dataset it was trained on (e.g., OpenAI-WIT-400M or DFN’s high-quality dataset). These strategies should always be orthogonal to the use of any newly trained CLIP-style models like DFN and might also be compatible with methods using external models like FAST and BLIP-2.
Our Contributions. We introduce two innovative approaches for data quality characterization in multi-modal contexts: negCLIPLoss and Norm Similarity Score (NormSim). negCLIPLoss addresses the limitations of traditional CLIPScores by normalizing scores using the similarity between a sample and its contrastive pairs, effectively reducing biases such as those introduced by generic terms like "image". This method applies universally across various embedding models (see Fig. 2). On the other hand, NormSim leverages available knowledge about downstream tasks to enhance data filtering. It computes the vision similarity between a training sample and a target dataset using a $p$-norm, where significant parameters are typically $p=2$ or $\infty$. Both metrics, negCLIPLoss and NormSim, complement each other in optimizing data selection for specific tasks (see Fig. 3).
To illustrate the effectiveness of our methods, we use a widely used benchmark DataComp (gadre2023datacomp). We show that, by simply replacing CLIPScores with negCLIPLoss and utilizing NormSim, we can exceed the best OAI-CLIP(L/14)-based baseline by 5.3% on ImageNet-1k and 2.8% on average across 38 downstream tasks, which is similar or even better than the performance achieved by many external-resources-based methods. Moreover, by merging the subset selected by negCLIPLoss and NormSim with the subset selected by current state-of-the-art method “HYPE $\cup$ DFN”, we can further improve it by 0.9% on both ImageNet-1k and on average 38 downstream tasks. We can also achieve a 0.8% improvement on average 38 tasks over "HYPE $\cup$ DFN" using only the data selected by DFN and our strategies. Besides, compared to existing methods, our approach saves a significant amount of computational time on data reprocessing as shown in Table 4.
2 Problem Setup
Data Filtering on Multimodal Dataset. We are given a training dataset $D_{\text{train}}=\{x^{v},x^{l}\}$, where $(x^{v},x^{l})\in\mathbb{R}^{D}$ is the image-text (vision-language) training pair. Our goal is to identify a subset $S\subset D_{\text{train}}$ that maximizes the zero-shot accuracy of the CLIP model on some downstream tasks when $S$ is used to train the CLIP model. CLIP. We denote the teacher CLIP model as $\bar{f}_{vl}$. For any pair $x^{vl}$, the model outputs a normalized unit-vector $\bar{f}_{vl}(x^{vl})$. If $X^{vl}:=\{x_{1}^{vl},\ldots,x_{m}^{vl}\}$ denotes a dataset containing $m$ samples, then we define $\bar{f}_{vl}(X^{vl})=[\bar{f}_{vl}(x_{1}^{vl}),\ldots,\bar{f}_{vl}(x_{m}^{vl})]^{\top}\in\mathbb{R}^{m\times d}$. The popular filtering metric “CLIPScore” is defined as $\langle\bar{f}_{v}(x^{v}),\bar{f}_{l}(x^{l})\rangle\in[-1,1]$. Dataset and model. We follow the pipeline of Datacomp (gadre2023datacomp) to standardize the training and evaluation process as previous papers  (nguyen2023improving; maharana2023d2; maini2023t; fang2023data; mahmoud2023sieve; bai2023qwen).More in Sec. 4.
Figure 1: Illustration of negCLIPLoss. CLIPScore may underestimate (bottom left) or overestimate (bottom right) the quality of image-text pairs. However, this issue can be mitigated by simply adding a normalization term like negCLIPLoss. “Top X%” denotes that the score represents the top Xhigh values within the entire dataset.
3 Data Filtering Strategy
3.1 negCLIPLoss: A Better Metric than CLIPScore
In this section, we introduce a better and statistically interpretable quality metric called negCLIPLoss, which directly replaces the common metric CLIPScore. Fig. 1 illustrates how negCLIPLoss works. As the name suggested, this metric is inspired by the standard CLIP loss used in the actual training process of the teacher CLIP model:
$\small\ell_{B^{*}}(x_{i}^{vl})=-\frac{1}{2}\left[\log\frac{\text{exp}(\bar{f}_{v}(x_{i}^{v})^{\top}\bar{f}_{l}(x_{i}^{l})/\tau)}{\sum_{j\in B^{*}}\text{exp}(\bar{f}_{v}(x_{i}^{v})^{\top}\bar{f}_{l}(x_{j}^{l})/\tau)}+\log\frac{\text{exp}(\bar{f}_{v}(x_{i}^{v})^{\top}\bar{f}_{l}(x_{i}^{l}))/\tau}{\sum_{j\in B^{*}}\text{exp}(\bar{f}_{v}(x_{j}^{v})^{\top}\bar{f}_{l}(x_{i}^{l})/\tau)}\right]$ (1)
Here $B^{*}$ is the random batch where $i$-th sample belongs during the final training step, and $\tau$ is the learnable temperate parameter. Notably, the teacher loss differs from CLIPScore primarily by a normalization term $\mathcal{R}^{*}$ as follows:
$\small-\tau\cdot\ell_{B^{*}}(x_{i}^{vl})=\underbrace{\bar{f}_{v}(x_{i}^{v})^{\top}\bar{f}_{l}(x_{i}^{l})}_{\text{CLIPScore}}-\underbrace{\frac{\tau}{2}\left[\log\sum_{j\in B^{*}}\exp(\frac{\bar{f}_{v}(x_{i}^{v})^{\top}\bar{f}_{l}(x_{j}^{l})}{\tau})+\log\sum_{j\in B^{*}}\exp(\frac{\bar{f}_{v}(x_{j}^{v})^{\top}\bar{f}_{l}(x_{i}^{l})}{\tau})\right]}_{\text{normalization term }\mathcal{R}^{*}}$
In practice, since the training dataset of teacher CLIP models, like OAI-WIT400M (radford2021learning), and the actual batch divisions $B^{*}$ is inaccessible, we randomly select $K$ batches from the student model’s training data and use the averaged results from $\{B_{k}\}_{i=1}^{K}$ to estimate the normalization term $\mathcal{R}^{*}$ on $B^{*}$, as shown in Eqn. 2:
$\small\text{negCLIPLoss}(x_{i}^{vl}):=-\frac{\tau}{K}\sum_{k=1}^{K}\ell_{B_{k}}(x_{i}^{vl})\ \approx\ \text{CLIPScore}(x_{i}^{vl})-\mathcal{R}^{*}$ (2)
Figure 2: Comparison of negCLIPLoss and CLIPScore across different downsampling ratios on DataComp-medium.
Here $\{B_{k}\}_{i=1}^{K}$ are some batches randomly selected from the student model’s training data and $x_{i}\in B_{k},\forall k$. We choose $K=10$ in our experiments, but any sample size larger than 5 is sufficiently stable (Appendix E.1). In Appendix D.2 we also show that the computational cost introduced by $\mathcal{R}$ remains negligible compared to other baselines. Temperature $\tau$ and batch size $|B^{*}|$ can be directly obtained from the pretrained teacher model, and the algorithm process of calculating negCLIPLoss is in Appendix D.3.
Motivation behind negCLIPLoss. Existing works like LESS (xia2024less) in NLP, CoDis (Xia_2023_combatting) in CV, and RHO (pmlr-v162-mindermann22a-prioritized) in data scheduling utilize loss-guided data selection. Despite that, the suitability of this approach for multi-modal contrastive learning remains uncertain. We address this by demonstrating that negCLIPLoss consistently performs as well as or better than CLIPScore (see Fig. 2). The reason is that negCLIPLoss incorporates a normalization term that corrects the biases inherent in CLIPScore (Fig. 1). This term adjusts for cases where text or image embeddings match with multiple contrastive pairs, not just their intended counterparts. For instance, terms like "Image" or "Photo" may result in falsely high CLIPScores as they can match various visual contents. Similarly, simple visual elements can align with numerous general descriptors, lowering their relative negCLIPLoss despite high initial CLIPScores. In contrast, specific words, such as "Islands Harbor" and "American football", have fewer potential matches, resulting in a higher relative negCLIPLoss. This illustrates how negCLIPLoss provides a more accurate measure in multi-modal contexts.
[…]
3.2 NormSim: A New Training-Target Similarity Metric
Our proposed negCLIPLoss is a universal approach to improve filtering performance by estimating quality better, and it does not rely on any downstream task. Now, if we can access some knowledge of the downstream tasks, we could further improve the performance by using a vision-only $p$-norm similarity to target data metric to measure the relationship between each training sample and the downstream target data. We will discuss the reason to use vision-only embedding later in this section.
Specifically, we assume access to the target set of downstream tasks and denote them as $X_{\text{target}}=\{x_{\text{target},(1)},\ldots,x_{\text{target},(m)}\}$, where each $x_{\text{target},(i)}\in\mathbb{R}^{d}$ is i.i.d.-sampled from the target downstream distribution $\mathcal{P}_{\text{target}}$ [^1] Although out-of-distribution tasks like “WILDS” have distribution shift between training data and test data, they still provides useful information of the test data., but without overlapping with the test set. Then, for each training sample $x^{vl}$ and the corresponding target set $X_{\text{target}}$, NormSim is defined as:
$\small\text{{NormSim}}_{p}(X_{\text{target}},x):=\|\bar{f}_{v}(X_{\text{target}}^{v})\bar{f}_{v}(x^{v})\|_{p}$ (3)
We select the subset $S$ of size $N$ by greedily choosing the samples with the largest NormSim scores. We consider $p=\infty$ in the main paper and left $p=2$ in Appendix A.1. When $p=\infty$, a training sample will be selected if it has enough high similarity to any target sample. Note that this is different from the nearest-neighbor-based method used in image-based filtering (gadre2023datacomp), where they are trying to find the nearest training sample of every target sample. In detail, $\text{NormSim}_{\infty}$ can be regarded as:
$\small S=\arg\max_{|S|=N}\sum_{i\in S}\max_{x_{t}\in X_{\text{target}}}\bar{f}_{v}(x_{t}^{v})^{\top}\bar{f}_{v}(x^{v}_{i})$ (4)
In Appendix E.4, we also show that our $\text{NormSim}_{\infty}$ can outperform the nearest neighbor selection on the downstream target tasks. Here, we show an example selected via the $\text{NormSim}_{2}$(ImageNet-1k) in Fig. 3 (Refer to Appendix F for more visualization.), showing that this vision-target-aware method is complementary to the quality-based one. Please refer to Appendix A.1 for more discussions.
4 Experimental Results
In this section, we evaluate the performance of negCLIPLoss and NormSim, aiming to address the following questions: Q1: Given a fixed CLIP teacher model, can our methods more effectively utilize CLIP embeddings for data filtering? Q2: Are our methods applicable to diverse CLIP teacher models with varying architectures or different pretrained datasets? Q3: How does our method compare to other leading approaches that utilize external models or multimodal datasets? Could our method be compatible with these methods?
4.1 Setup
We adhere to the standardized training and evaluation protocols of the DataComp benchmark (gadre2023datacomp). We employ the medium-scale training configuration of DataComp (DataComp-medium). It provides a substantial dataset comprising 128 million low-quality, web-curated image-text pairs to be filtered. Once the data subset is obtained by some data filtering strategy, it will be used to train a fixed CLIP-B/32 model in a fixed training budget that allows the model to pass 128 million data points an epoch. Therefore, smaller subsets will be repeated more frequently, ensuring a fair comparison. We note that the size of the DataComp dataset becomes smaller over time since some URLs of images become invalid [^2] See https://github.com/mlfoundations/datacomp/issues/3 and similar issues in maharana2023d2. , and we only successfully downloaded about 110M data. Therefore, the results of baselines on the leaderboard do not apply to our datasets. So we reproduce top baselines on the leaderboard with their public UIDs of the selected data [^3] The results of MLM (wang2024finetuned) are from their paper.. More details about evaluation tasks and architecture are in Appendix D.1.
4.2 Baselines
We restate the three current research directions mentioned before based on how much external resources are employed: (D1) using OAI CLIP alone while optimizing embedding employment strategies, (D2) training and using a more advanced CLIP embedding model based on external data, and (D3) utilizing non-CLIP external models to aid data selection [^4] Note that D2 and D3 may also incorporate strategies from D1. For example, CLIPScore (D1) has been used in almost all the top methods. Therefore, we categorize baselines by the largest possible category they encompass.. In our experiment parts, we consider baselines following the above categorization: (1) D1, including CLIPScore (hessel2021clipscore), Image-based filtering (gadre2023datacomp), $\mathbb{D}^{2}$ Pruning (maharana2023d2). (2) D2, D3, including DFN (fang2023data), HYPE (kim2024hype), HYPE $\cup$ DFN (previous state-of-the-art), T-MARS (maini2023t), Devils (yu2023devil), and MLM (wang2024finetuned). Please refer to Fig. 4, Appendix C, and Appendix D.5 for more details. Intuitively, performance should be ranked as D2, D3 > D1. However, our results show that cross-setting comparisons are possible and our D1 methods can perform similar or even better than most of D3 methods.
Table 1: Results on DataComp-medium from methods that use only OpenAI’s CLIP-L/14 model, i.e., all methods are from the D1 category. The “dataset size” represents the size of the subset obtained from different approaches. NormSim(IN-1k) denotes using the training data of ImageNet-1k as the target while NormSim(Target) represents using that of all 24 available downstream tasks. Complete Results in Appendix E.2.
Filtering Strategy Dataset IN-1k IN Shift VTAB Retr. Avg.
Size (1 task) (5) (11) (3) (38)
No filtering (gadre2023datacomp) 110M 17.3 15.0 25.2 21.3 25.6
CLIPScore (20%) (hessel2021clipscore) 22M 25.4 22.7 31.8 22.0 31.0
CLIPScore (30%) (hessel2021clipscore) 33M 26.4 23.6 32.6 24.5 32.2
Image-based (gadre2023datacomp) 24M 25.5 21.9 30.4 24.6 29.9
CLIPScore (30%) $\cap$ Image-based 11M 27.4 23.9 31.9 21.4 30.8
$\mathbb{D}^{2}$ Pruning (maharana2023d2) 22M 23.2 20.4 31.4 18.7 29.5
negCLIPLoss (20%) 22M 27.4 23.8 33.7 23.7 32.5
negCLIPLoss (30%) 33M 27.9 24.6 33.2 25.1 32.9
CLIPScore (30%) $\cap$ $\text{NormSim}_{\infty}$(IN-1k) 22M 29.7 25.9 33.7 24.1 33.7
CLIPScore (30%) $\cap$ $\text{NormSim}_{\infty}$(Target) 22M 30.2 26.2 35.0 23.4 33.9
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(IN-1k) 22M 31.9 27.3 34.8 25.0 34.4
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target) 22M 31.7 27.2 36.0 26.0 35.0
4.3 Main Results and Discussions
4.3.1 D1 Category (Q1)
In Table 1, we compare D1 methods where only the OAI CLIP model is allowed to be used. We show that our Methods leverage OAI CLIP-L/14 better. First, negCLIPLoss outperforms CLIPScore on all metrics, regardless of whether it is used alone or combined with other methods. These results support our claim that negCLIPLoss can more accurately estimate the data quality. Second, when target knowledge is available, $\text{NormSim}_{\infty}$ can further improve filtering significantly. Especially, compared with the best baseline “CLIPScore (30%)”, our “negCLIPLoss $\cap$ $\text{NormSim}_{\infty}$(Target)” improves 5.3% on ImageNet-1k and 2.8% on average 38 downstream tasks, respectively. Later in Table 3 we will see that this result outperforms all the D3 baselines except DFN $\cup$ HYPE. And we can also see that using ImageNet-1k-only target data can still produce good results.
Table 2: Results on DataComp-medium from the top methods that use only OpenAI’s CLIP-B/32 model or public version of DFN (DFN-P).
Strategy Size IN-1k VTAB Avg.
OAI CLIP-B/32
CLIPScore (20%) 22M 27.0 33.0 32.2
CLIPScore (30%) 33M 27.6 33.6 33.2
negCLIPLoss (20%) 22M 28.9 34.3 33.0
negCLIPLoss (30%) 33M 28.8 33.7 33.6
negCLIPLoss (30%) 22M 32.4 35.9 35.2
$\cap$ $\text{NormSim}_{\infty}$(Target)
DFN-P
CLIPScore (15%) 16M 25.9 32.9 31.6
CLIPScore (17.5%) 19M 30.2 34.1 33.8
CLIPScore (20%) 22M 29.7 33.0 33.1
CLIPScore (30%) 33M 28.4 33.2 32.7
negCLIPLoss (15%) 16M 31.3 35.8 34.6
negCLIPLoss (17.5%) 19M 31.2 35.7 34.7
negCLIPLoss (20%) 22M 30.7 33.6 33.8
negCLIPLoss (30%) 33M 28.9 33.4 33.2
negCLIPLoss (30%) 22M 29.4 33.5 32.5
$\cap$ $\text{NormSim}_{\infty}$(Target)
negCLIPLoss (17.5%) 16M 31.6 37.2 35.7
$\cap$ $\text{NormSim}_{\infty}^{\text{B/32}}$(Target)
4.3.2 Universality (Q2)
To evaluate whether our method applies to other CLIP teacher models, we replaced OAI-L/14 with OAI-B/32 and DFN-P as embedding models. Results are in Table 2 and Appendix E.3.555The original DFN paper selects a subset comprising 19.2M data points, which accounts for approximately $17.5\%$ of our dataset and $15\%$ of their dataset, we incorporate these sampling ratios into our comparison. We conclude that (1) negCLIPLoss can be applied to different CLIP embedding models. Our proposed negCLIPLoss can consistently outperform CLIPScore on all downstream tasks for different filtering ratios and models, like a 0.5%-5.4% increase on ImageNet-1k. (2) Embedding required by NormSim should have good downstream performance. When combining negCLIPLoss with $\text{NormSim}_{\infty}$, OAI-B/32 obtains results even better than those using OAI-L/14 (Table 1), while DFN-P achieves results even worse than using negCLIPLoss alone. The reason is that, unlike OAI CLIP, DFN-P is specially designed for data filtering at the expense of downstream task performance [^6] For example, the ImageNet-1k accuracy for DFN-P, OAI CLIP-B/32, and OAI CLIP-L/14 are 45%, 63%, and 75%, respectively.. This indicates that the embeddings obtained from DFN on target data might be highly unreliable, leading to inaccurate NormSim. Furthermore, if we use DFN-P for negCLIPLoss but OAI-B/32 for NormSim ("negCLIPLoss (17.5%) $\cap$ $\text{NormSim}_{\infty}^{\text{B/32}}$(Target)"), we can achieve the best results.
4.3.3 Comparison with D2 & D3 Categories (Q3)
In this part, we compare all the D2 & D3 baselines mentioned in Sec. 4.2 together with our best strategy in Table 3.  [^7] For “A $\cup$ B” mentioned in Table 3, we follow the way of Kim et al. (kim2024hype) to merge the data, which generates the subset separately for each method and then merge them. This will result in oversampling the shared data, which is intuitively more important. We claim: (1) Our methods can outperform most of the D3 methods. In Table 3, we show that without using any external models or data, our best combination, i.e., using OAI-B/32 for “negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target)” (Ours (20%)), still outperforms all methods except DFN and “DFN $\cup$ HYPE”. This answers the first part of Q3 and further indicates that some external models may be redundant since CLIP embeddings already contain necessary information. (2) We can further improve the SOTA method. We can further boost the performance of the current SOTA method “HYPE $\cup$ DFN” by 0.9% on both ImageNet-1k and on average 38 downstream tasks, and close results can be achieved even without combining HYPE which utilizes the external embedding model MERU (desai2023hyperbolic). Here we use the data selected by both OAI-B/32 and L/14, which we found is more robust than using one of them alone. Our better results answer the second part of Q3, that is, our methods can be compatible with other D2&D3 methods.
Table 3: Results of all D1&D2&D3 top methods on DataComp-medium. “Ours (20%)” refers to use “negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target)” to get 20% of original data, while “Ours (10%)” denotes applying “negCLIPLoss (20%) $\cap$ $\text{NormSim}_{\infty}$(Target)” to get 10%. And we use “*” to indicate the case where we choose the intersection of the data selected by using OAI CLIP-B/32 and OAI CLIP-L/14 separately, which results in about 15M data for “Ours (20%)*” and 7.4M data for “Ours (10%)*”.
Type Filtering Strategy Size IN-1k IN Shift VTAB Retr. Avg.
(1) (5) (11) (3) (38)
D2 DFN (fang2023data) 16M 36.0 30.1 36.2 27.0 35.4
D3 T-MARS (maini2023t) 22M 30.8 26.3 34.8 25.4 34.1
D3 Devil (yu2023devil) 20M 31.0 26.7 35.9 24.7 34.5
D3 MLM (wang2024finetuned) 38M 30.3 25.6 36.0 29.0 34.5
D3 HYPE (kim2024hype) 10M 30.3 25.8 34.3 22.2 31.9
D3 DFN $\cup$ HYPE (kim2024hype) 20M 36.4 30.8 38.5 28.0 36.8
D1 Ours (20%) 22M 32.4 27.4 35.9 26.3 35.2
D3 DFN $\cup$ Ours (20%)* 23M 36.4 30.9 38.6 28.1 37.6
D3 DFN $\cup$ HYPE $\cup$ Ours (10%)* 22M 37.3 31.4 38.5 27.6 37.7
5 Conclusion and Reproducibility Statement
In this paper, we introduce negCLIPLoss and NormSim to enhance data selection in multimodal contrastive learning without relying on external resources. negCLIPLoss provides a more accurate quality metric compared to the commonly used CLIPScore, while NormSim measures the similarity between pretraining data and target data for known downstream tasks. Experiments show that our method demonstrates both effectiveness (Table 1-3) and efficiency (Table 4). Additionally, negCLIPLoss and NormSim are compatible with existing top techniques, allowing us to achieve a new state-of-the-art by combining them. All experiments are reproducible using code https://github.com/ypwang61/VAS/tree/negcliploss_normsim.
Appendix A Additional Discussion
A.1 Discussion about NormSim
Necessity of using vision-only information.
We use only the visual information $x^{v}$ instead of multi-modal information $x^{vl}$ for measuring similarity. This is because common crawled text often has brief captions, making the OAI CLIP language embedding weaker than its visual embedding model (gadre2023datacomp; shen2021much; zeng2022multi; yamada2022lemons). Consequently, the language part cannot characterize the pre-training and downstream task distribution as well as the visual part. This phenomenon is also observed in Gadre et al. (gadre2023datacomp), where image-based filtering (select data whose image embeddings are similar to that from ImageNet-1k) outperforms text-based filtering (select data whose captions contain words from ImageNet-21k). More ablation studies are provided in Appendix E.5.
Using proxy when downstream $X_{\text{target}}$ is inaccessible.
Surprisingly, we show that the 2-norm can also be used when only the pre-training set is available. In this case, we construct a proxy “target” set from the pre-training set itself. Specifically, let $S_{i}$ be the selected subset at step $i$, then we treat the current $S_{i}$ as the proxy “target” set. To construct the next smaller set, we greedily select the next data batch $S_{i+1}$ satisfying $\operatorname*{arg\,max}_{S_{i+1}\subset S_{i}}\sum_{x\in S}\text{{$\text{NormSim}_{2}$}}(S_{i},x),$ until reaching an N size subset. We call this approach $\text{{NormSim}}_{2}$-D (Dynamic) and will specify the algorithm details in Appendix D.4.
Theoretical justification.
Unlike many existing methods that force diversity by selecting training samples around each ${\bm{x}}_{\text{target}}$, our strategy greedily maximizes similarity without considering actual coverage. For the $p=2$ case, we demonstrate that maximizing $\text{NormSim}_{2}$ is optimal under a linear model $\bar{f}_{v}$, as shown in Appendix B. Our theorem also provides error guarantees for noisy embeddings and explains when vision-only embeddings outperform combined vision and language embeddings. Recent work by Joshi et al. (pmlr-v238-joshi24a) provides a similar analysis but focuses on high-quality data and cross-variance between images and texts. This approach is less effective than image-only methods for filtering noisy datasets, as discussed above.
Appendix B Theoretical Interpretation
In this section, we give a theoretical justification on the NormSim metric when $p=2$ under the linear model assumptions when low quality image and mismatched text has already been removed. In other words, we mainly focus on the following strategy.
$\small S=\arg\max_{|S|=N}\sum_{i\in S}\bar{f}_{v}(x_{i}^{v})^{\top}\underbrace{\left(\frac{1}{|X_{\text{target}}|}\sum_{x_{t}\in X_{\text{target}}}\bar{f}_{v}(x_{t}^{v})\bar{f}_{v}(x_{t}^{v})^{\top}\right)}_{\bar{\Sigma}_{\text{target\_proxy}}}\bar{f}_{v}(x_{i}^{v})$ (5)
B.1 Theoretical Setup
Training data.
For any ${\bm{x}}^{v},{\bm{x}}^{l}\in\mathbb{R}^{d}$ observable image and text training pairs, we define ${\bm{z}}^{v},{\bm{z}}^{l}$ to be the corresponding latent vectors which contain all semantically pertinent information about our tasks of interest. Similar to previous theoretical work (nakada2023understanding), we assume each i.i.d pair ${\bm{z}}^{vl}$ follows zero-mean sub-gaussian distribution whose cross-covariance satisfies
$\text{Cov}({\bm{z}}^{v},{\bm{z}}^{l})=\Sigma_{\text{train}}=\text{diag}(\sigma_{1},\sigma_{2},\ldots),$ $\|{\bm{z}}^{vl}\|=1$
and each ${\bm{x}}^{vl}$ is generated based on a linear model such that
${\bm{x}}^{vl}=G^{*}_{vl}{\bm{z}}^{vl}+{\bm{\xi}}^{vl}.$
Here $G^{*}_{vl}\in O_{d\times r}$ is the othonormal ground truth representation mapping from the latent vector space to the input space, and $\xi^{vl}\sim\mathcal{N}(0,I_{d})$ are i.i.d. random noise.
Also we denote the cross covariance of any finite dataset $S^{\prime}$ (e.g. the given train set $D_{\text{train}}$) as $\Sigma_{S^{\prime}}$.
Test data.
For any zero-shot downstream task, we assume it shares almost same data generation process as the training set, except its the cross-covariance $\Sigma_{\text{target}}$ does not necessarily equal $\Sigma_{\text{train}}$, which necessitate the choice of $\bar{\Sigma}_{\text{target\_proxy}}$.
CLIP embedding model as teacher.
Under the linear model assumption, we have a teacher model $\bar{f}_{vl}=\bar{G}_{vl}$, whose generated clip embedding can partially recover the ground truth hidden vector ${\bm{z}}^{vl}$ with error.
Formally, we say teacher has $\epsilon_{v}^{n}$ error if for all possible $n$ budget subsets $S\subset D_{\text{train}}$,
$\frac{1}{|S|}\left\|\sum_{{\bm{x}}^{vl}\in S}\bar{G}_{v}^{\top}{\bm{x}}^{v}({\bm{x}}^{v})^{\top}\bar{G}_{v}-\sum_{{\bm{x}}^{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{v})^{\top}\right\|_{*}\leq\epsilon_{v}^{n}$
where the same notation applies for the language modal. By the orthonormal assumption on the ground truth matrix $G_{vl}^{*}$, we see that $\bar{G}_{v}^{\top}$ is aiming to inverting the map. In addition, we say the teacher has $\epsilon_{v*l}^{n}$ cross modal error
$\frac{1}{|S|}\|\sum_{{\bm{x}}^{vl}\in S}\bar{G}_{v}^{\top}{\bm{x}}^{v}({\bm{x}}^{l})^{\top}\bar{G}_{l}-\sum_{{\bm{x}}^{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{l})^{\top}\|_{*}\leq\epsilon_{v*l}^{n}$
When all $\epsilon_{v}^{n},\epsilon_{l}^{n},\epsilon_{v*l}^{n}\to 0$ as $n\rightarrow\infty$, then we say the teacher is strong for both modalities. But it might also be possible that only one modal, for example, visual is strong. That is $\epsilon_{v}^{n}\to 0,\epsilon_{l}^{n},\epsilon_{v*l}^{n}\gg\epsilon_{v}^{n}$.
Model and training.
According to Lemma 4.1 in (nakada2023understanding), using the CLIP loss to optimize the linear model has approximately the same training dynamics as using the regularized linear loss. Therefore, here we assume that we are learning $G_{v},G_{l}$ by maximizing the clip score gap between the contrastive pairs, plus a regularizer,
$\min_{G_{v},G_{l}}{\mathcal{L}}_{S}^{\rho}(G_{v},G_{l}):=\min_{G_{v},G_{l}}\frac{\sum_{i\in S}\sum_{j\in S}(s_{ij}-s_{ii})}{|S|(|S|-1)}+\frac{\rho}{2}\frac{|S|}{|S|-1}\|G_{v}G_{l}^{\top}\|_{F}^{2}$
where $s_{ij}:=\langle G_{v}^{\top}{\bm{x}}^{v}_{i},G_{l}^{\top}{\bm{x}}^{l}_{j}\rangle$ and $\rho>0$ is some regularizer-related constant. Note that this objective maximizes self-similarity and minimizes similarity between disparate pairs. Note that this “loss” can be negative, avoiding the trivial null solution of all zeros. We denote this training process from any given $S$ as $G_{vl}={\mathcal{A}}^{\rho}(S)$.
Goal and metric.
Under the same principle as our training loss function, we measure the performance of any learnt $G_{v},G_{l}$ on some downstream task with distribution ${\mathcal{D}}_{\text{target}}$ as test loss ${\mathcal{L}}_{\text{target}}(G_{v},G_{l}):=$
$\mathbb{E}_{\begin{subarray}{c}{\bm{x}}^{vl}\sim{\mathcal{D}}_{\text{target}}\\ {\bm{x}}^{vl}_{2}\sim{\mathcal{D}}_{\text{target}}\end{subarray}}(\langle G_{v}^{\top}{\bm{x}}^{v},G_{l}^{\top}{\bm{x}}_{2}^{l}\rangle-\langle G_{v}^{\top}{\bm{x}}^{v},G_{l}^{\top}{\bm{x}}^{l}\rangle)$
This is inspired by the following classification accuracy. Assume that the test data including $C$ class, and the class distribution is $\mathcal{C}$. For every class $c$, the training data ${\bm{x}}=({\bm{x}}^{v},{\bm{x}}^{l})$ satisfies distribution $\mathcal{P}_{c}$. We further assume the corresponding classification templates are $\{{\bm{x}}_{c}\}_{c=1}^{C}$. Thus we define classification accuracy as
$\text{AC}(G_{v},G_{l})=\mathbb{E}_{c,c^{\prime}\sim\mathcal{C}\times\mathcal{C}}\left[\mathbb{E}_{{\bm{x}}_{i}\sim\mathcal{P}_{c}}\mathbf{1}[s_{ic}>s_{ic^{\prime}}]\right]$
Therefore our goal is to minimize its gap between the best hind-side subset, for any $\rho$, without budget constraints,
$\Delta^{\rho}(S)={\mathcal{L}}_{\text{target}}(\hat{G}_{vl})-\min_{S^{\prime}\in D_{\text{train}}}{\mathcal{L}}_{\text{target}}({\mathcal{A}}^{\rho}(S^{\prime})),\hat{G}_{vl}={\mathcal{A}}^{\rho}(S)$
B.2 Generalization Guarantees
We now provide theoretical guarantees and postpone our proof into Appendix B.3. Firstly, we are going to prove the intuition behind $\text{NormSim}_{2}$score.
Lemma B.1 (Intuition behind $\text{NormSim}_{2}$).
With high probability at least $1-\frac{1}{|S|d}$, suppose the hind-side best subset has at least $\underline{n}$ number of samples, then we have
$\Delta^{\rho}(S)$ $=\underbrace{\frac{1}{\rho}\max_{S^{\prime}\in D_{\text{train}}}\left(\operatorname{Tr}\left(\Sigma_{\text{target}}(\Sigma_{S^{\prime}}-\Sigma_{S})\right)\right)}_{\text{{$\text{NormSim}_{2}$}related term}}+\quad\underbrace{{\mathcal{O}}\left(\sqrt{\frac{d\log(d|S|)}{\underline{n}}}+\sqrt{\frac{d\log(d|S|)}{|S|}}\right)}_{\text{noise}}$
Proof sketch.
❶ Under the assumption that both ${\bm{z}}^{vl},\xi_{vl}$ is zero-mean, maximizing the clip score gap is equivalent to maximizing the clip score of the same sample.
${\mathcal{L}}_{\text{target}}(\hat{G}_{v},\hat{G}_{l}):=-\mathbb{E}_{{\bm{x}}^{vl}\sim{\bm{{\mathcal{D}}}}_{\text{target}}}\langle\hat{G}_{v}^{\top}{\bm{x}}^{v},\hat{G}_{l}^{\top}{\bm{x}}^{l}\rangle$
❷ By minimizing the regularized training loss ${\mathcal{L}}_{S}^{\rho}(G_{v},G_{l})$ using Eckart-Young-Mirsky Theorem, we get a closed form solution of $\hat{G}$ as
$\hat{G}_{v}\hat{G}_{l}^{\top}\approx\frac{1}{\rho}G_{v}^{*}\Sigma_{S}\cdot(G_{l}^{*})^{\top}+\text{noise depend on $S$}$
❸ Combining the result in ❷ and ❶, we have
${\mathcal{L}}_{\text{target}}(\hat{G}_{vl})\approx-\frac{1}{\rho}\operatorname{Tr}\left(\Sigma_{\text{target}}\Sigma_{S}\right)-\text{noise depend on $S$}$
The same analysis can be applied on $\min_{S^{\prime}\in D_{\text{train}}}{\mathcal{L}}_{\text{target}}({\mathcal{A}}(S^{\prime}))$ as well. Rearranging these two equations gives us the final result. ∎
This lemma shows the the $\Delta(S)$ is depend on the $\text{NormSim}_{2}$-related term and the noise term which comes from $\xi$. When $\underline{n}$ and $|S|$ is large enough, then the $\text{NormSim}_{2}$-related term will become dominant. This aligns with our practice experience that the final performance is less sensitive to the small variation in the number of select data as long as that is sufficient. Moreover, in some special cases where test distribution has identity cross-variance, then sampling by greedily choosing CLIP score might be enough.
Now we are ready to give a proof on the choice of $\bar{\Sigma}_{\text{target}}$ and visual-only information. Specifically, the strategy error mainly comes from (1). The unknown test distribution shift from training. (2). The unobservable ground truth $\Sigma_{S}$. To tackle error (1), we assume some prior knowledge on test by using the proxy test variance $\bar{\Sigma}_{\text{target}}$. To tackle the error (2), there are two possible solutions as shown below. Based on the theoretical interpretation, we should choose different strategy based on the property of the teacher embedding model.
$S_{\text{vision+language}}=\operatorname*{arg\,max}_{S}\operatorname{Tr}\left(\bar{\Sigma}_{\text{target}}(\sum_{{\bm{x}}^{vl}\in S}\bar{G}_{v}^{\top}{\bm{x}}^{v}({\bm{x}}^{l})^{\top}\bar{G}_{l})\right)$
$S_{\text{vision only}}=\operatorname*{arg\,max}_{S}\operatorname{Tr}\left(\bar{\Sigma}_{\text{target}}(\sum_{{\bm{x}}^{vl}\in S}\bar{G}_{v}^{\top}{\bm{x}}^{v}({\bm{x}}^{v})^{\top}\bar{G}_{v})\right)$
Theorem B.1 (Main).
Under the assumption of Lemma B.1,
$\Delta^{\rho}(S)$ $\leq\text{noise}+\frac{1}{\rho}\|\bar{\Sigma}_{\text{target}}-\Sigma_{\text{target}}\|\|\Sigma_{S}-\Sigma_{\text{best}}\|_{*}$
$+\frac{1}{\rho}\begin{cases}\epsilon_{v*l}^{S}\quad\text{(vision+language)}\\ \epsilon_{v}^{S}+\sqrt{1-\frac{1}{|S|}\sum_{i\in[S]}\langle{\bm{z}}^{v},{\bm{z}}^{l}\rangle)}\quad\text{(vision only)}\end{cases}$
Firstly, it is evident that the greater the difference between $\bar{\Sigma}_{\text{target}}$ and $\Sigma_{\text{target}}$, the less improvement we can expect. Moreover, in scenarios where $\epsilon_{l}$ is large (indicating lower accuracy in the language part) while $\epsilon_{v}$ is small (indicating higher accuracy in the vision part), it may be advisable to opt for vision-only embeddings. However, the learner should also consider the term $\sqrt{1-\frac{1}{|S|}\sum_{i\in S}\langle{\bm{z}}^{v},{\bm{z}}^{l}\rangle}$, which represents the alignment between the ground truth visual and language latent vectors, essentially reflecting the intrinsic quality of the data. If this term is already significant, relying solely on vision information as a proxy for language information could lead to suboptimal results.
B.3 Detailed proofs
Lemma B.2.
Let
$\hat{G}_{v},\hat{G}_{l}=\arg\min_{G_{v},G_{l}\in\mathbb{R}^{d\times r}}\mathcal{L}(G_{v},G_{l})$ (6)
Then we have
$\hat{G}_{v}\hat{G}_{l}^{\top}=\frac{1}{\rho}G_{v}^{*}\Sigma_{S}(G_{l}^{*})^{\top}+P_{1}+P_{2}+P_{3}+P_{4}$ (7)
where noise terms $P_{i}$ are defined in (11) , (12), (13) and (14).
Proof.
Note that $s_{ij}=({\bm{x}}_{j}^{l})^{\top}G_{l}G_{v}^{\top}{\bm{x}}_{i}^{v}=\operatorname{Tr}(G_{v}^{\top}{\bm{x}}_{i}^{v}({\bm{x}}_{j}^{l})^{\top}G_{l})$, like the proof of Corollary B.1. in (nakada2023understanding), we have
$\small\mathcal{L}(G_{v},G_{l})$ $=$ $\frac{\sum_{i\in S}\sum_{j\in S}(s_{ij}-s_{ii})}{|S|(|S|-1)}+\frac{\rho}{2}\frac{|S|}{|S|-1}\|G_{v}G_{l}^{\top}\|_{F}^{2}$
$=$ $\frac{\sum_{i\in S}\sum_{j\in S}s_{ij}-|S|\sum_{i\in S}s_{ii}}{|S|(|S|-1)}+\frac{\rho}{2}\frac{|S|}{|S|-1}\|G_{v}G_{l}^{\top}\|_{F}^{2}$
$=$ $-\operatorname{Tr}\left(G_{v}^{\top}\left[\frac{1}{|S|-1}\sum_{i\in S}{\bm{x}}_{i}^{v}({\bm{x}}_{i}^{l})^{\top}-\frac{|S|}{|S|-1}\bar{\bm{x}}^{v}(\bar{\bm{x}}^{l})^{\top}\right]G_{l}\right)+\frac{\rho}{2}\frac{|S|}{|S|-1}\|G_{v}G_{l}^{\top}\|_{F}^{2}$
=:absent:=:= : $-\operatorname{Tr}(G_{v}^{\top}\Gamma G_{l})+\frac{\rho}{2}\frac{|S|}{|S|-1}\|G_{v}G_{l}^{\top}\|_{F}^{2}$
where $\bar{\bm{x}}^{vl}:=(\sum_{i\in S}{\bm{x}}_{i}^{vl})/|S|$. Then by the Eckart-Young-Mirsky Theorem (For example, Theorem 2.4.8 in Golub et al. (golub2013matrix)), we know that
$\arg\min_{G_{v}\in\mathbb{R}^{d\times r},G_{l}\in\mathbb{R}^{d\times r}}\mathcal{L}(G_{v},G_{l})$
$=$ $\arg\max_{G_{v}\in\mathbb{R}^{d\times r},G_{l}\in\mathbb{R}^{d\times r}}\operatorname{Tr}(G_{v}^{\top}\Gamma G_{l})-\frac{\rho}{2}\frac{|S|}{|S|-1}\|G_{v}G_{l}^{\top}\|_{F}^{2}$
[…]
where the notation $\mathrm{SVD}_{r}(\Gamma)$ means choosing the first $r$ components of the matrix $\Gamma$. Further note that
$\Gamma$ $=$ $\frac{1}{|S|-1}\sum_{i\in S}{\bm{x}}_{i}^{v}({\bm{x}}_{i}^{l})^{\top}-\frac{|S|}{|S|-1}\bar{\bm{x}}^{v}(\bar{\bm{x}}^{l})^{\top}$ (8)
=:absent:=:= : $P_{0}+P_{1}+P_{2}+P_{3}+P_{4}$ (9)
Here note that $\Sigma_{S}=\frac{1}{|S|}\sum_{i\in S}{\bm{z}}_{i}^{v}({\bm{z}}_{i}^{l})^{\top}$, we have $P_{i}$ as follows:
$P_{0}$ ::=:= $\frac{|S|}{|S|-1}G_{v}^{*}\cdot\Sigma_{S}\cdot(G_{l}^{*})^{\top}$ (10)
$P_{1}$ ::=:= $\frac{1}{|S|-1}G_{v}^{*}\sum_{i\in S}{\bm{z}}_{i}^{v}({\bm{\xi}}_{i}^{l})^{\top}$ (11)
$P_{2}$ ::=:= $\frac{1}{|S|-1}\sum_{i\in S}{\bm{\xi}}_{i}^{v}({\bm{z}}_{i}^{l})^{\top}(G_{l}^{*})^{\top}$ (12)
$P_{3}$ ::=:= $\frac{1}{|S|-1}\sum_{i\in S}{\bm{\xi}}_{i}^{(1)}({\bm{\xi}}_{i}^{(2)})^{\top}$ (13)
$P_{4}$ ::=:= $-\frac{|S|}{|S|-1}\bar{\bm{x}}^{v}(\bar{\bm{x}}^{l})^{\top}$ (14)
It’s clear that the rank of the matrix $P_{0}$ is no more than $r$, so $\mathrm{SVD}_{r}(P_{0})=P_{0}$. And for $i\in\{1,2,3,4\}$, $P_{i}$ are noise terms with $\mathbb{E}[P_{i}]=O$. ∎
Lemma B.3.
For any fixed $S$, w.h.p $1-\delta$ the noise term can be upper bounded by $\sqrt{\frac{d\log(1/\delta)}{|S|}}$
Proof.
To upper bound the P1 and P2, we have
$\|\sum_{i}{\bm{z}}_{i}^{vl}(\xi_{i}^{vl})^{\top}\|_{*}^{2}=\operatorname{Tr}\left(\sum_{i,j}\xi_{i}^{vl}({\bm{z}}_{i}^{vl})^{\top}{\bm{z}}_{j}^{vl}\xi_{j}^{vl}\right)=\sum_{i,j}({\bm{z}}_{i}^{vl})^{\top}{\bm{z}}_{j}^{vl}(\xi_{j}^{vl})^{\top}\xi_{i}^{vl}$
$\mathbb{E}\|\sum_{i}{\bm{z}}_{i}^{vl}(\xi_{i}^{vl})^{\top}\|_{*}^{2}=\mathbb{E}\left[\sum_{i}({\bm{z}}_{i}^{vl})^{\top}{\bm{z}}_{i}^{vl}(\xi_{i}^{vl})^{\top}\xi_{i}^{vl}\right]=|S|d$
Regarding each $({\bm{z}}_{i}^{vl})^{\top}{\bm{z}}_{j}^{vl}(\xi_{j}^{vl})^{\top}\xi_{i}^{vl}$ as weakly dependent variable, then by using Bernstein inequality, we have, with high probability $1-\delta$,
$\|\sum_{i}{\bm{z}}_{i}^{vl}(\xi_{i}^{vl})^{\top}\|_{*}^{2}\leq|S|d+\sqrt{d|S|^{2}\sigma_{\xi}^{2}\log(1/\delta)}\leq|S|d\sqrt{\log(1/\delta)}$
So $\frac{1}{|S|}\|\sum_{i}{\bm{z}}_{i}^{vl}(\xi_{i}^{vl})^{\top}\|_{*}\leq\sqrt{\frac{d\log(1/\delta)}{|S|}}$. Note that $\|\bar{\bm{x}}^{vl}\|\lesssim\sqrt{\frac{\log(|S|d)}{|S|}}$ (like Proposition 2.5 in Wainwright et al. (wainwright2019high)), it is easy to see that P3 ad P4 are the low order terms if $\delta\lesssim\frac{1}{|S|d}$. ∎
Lemma B.4 (Intuition behind VAS).
With high probability $1-\delta$, suppose the hind-side best subset has at least $\underline{n}$ number of samples, then we have
$\Delta(S)=\frac{1}{\rho}\max_{S^{\prime}\in D_{\text{train}}}\left(\operatorname{Tr}\left(\Sigma_{\text{target}}(\Sigma_{S^{\prime}}-\Sigma_{S})\right)\right)+\sqrt{\frac{d\log(1/\delta)}{\underline{n}}}+\sqrt{\frac{d\log(1/\delta)}{|S|}}$
Proof.
For any learnt $G_{v},G_{l}$ based on dataset $S$, we have
${\mathcal{L}}_{\text{test}}(G_{v},G_{l})$ $=\operatorname{Tr}(G_{v}^{\top}\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]G_{l})$
$=\operatorname{Tr}(\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]G_{l}G_{v}^{\top})$
$=\frac{1}{\rho}\operatorname{Tr}\left(\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]G_{l}^{*}\Sigma_{S}(G_{v}^{*})^{\top}\right)-\operatorname{Tr}\left(\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]\text{noise}_{S}\right)$
$=\frac{1}{\rho}\operatorname{Tr}\left((G_{v}^{*})^{\top}\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]G_{l}^{*}\Sigma_{S}\right)-\operatorname{Tr}\left(\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]\text{noise}_{S}\right)$
$=-\frac{1}{\rho}\operatorname{Tr}\left(\Sigma_{\text{target}}\Sigma_{S}\right)-\operatorname{Tr}\left(\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]\text{noise}_{S}\right)$
Here the first equation comes from Theorem B.3 and the third equation comes from Lemma B.2.
Consequently, we have
$-\min_{S^{\prime}\in D_{\text{train}}}{\mathcal{L}}_{\text{test}}({\mathcal{A}}(S^{\prime}))$ $=\max_{S^{\prime}\in D_{\text{train}}}\left(\frac{1}{\rho}\operatorname{Tr}\left(\Sigma_{\text{target}}\Sigma_{S^{\prime}}\right)+\operatorname{Tr}\left(\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]\text{noise}_{S^{\prime}}\right)\right)$
$\leq\frac{1}{\rho}\max_{S^{\prime}\in D_{\text{train}}}\left(\operatorname{Tr}\left(\Sigma_{\text{target}}\Sigma_{S^{\prime}}\right)\right)+\|\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}[{\bm{x}}^{v}({\bm{x}}^{l})^{\top}]\|\|\text{noise}_{S^{\prime}}\|_{*}$
$\leq\frac{1}{\rho}\max_{S^{\prime}\in D_{\text{train}}}\left(\operatorname{Tr}\left(\Sigma_{\text{target}}\Sigma_{S^{\prime}}\right)\right)+{\mathcal{O}}\left(\sqrt{\frac{d\log(1/\delta)}{\underline{n}}}\right)$
Therefore, we have the final result as
$\Delta(S)$ $={\mathcal{L}}_{\text{test}}(\hat{G}_{vl})-\min_{S^{\prime}\in D_{\text{train}}}{\mathcal{L}}_{\text{test}}({\mathcal{A}}(S^{\prime}))$
$=\frac{1}{\rho}\max_{S^{\prime}\in D_{\text{train}}}\left(\operatorname{Tr}\left(\Sigma_{\text{target}}(\Sigma_{S^{\prime}}-\Sigma_{S})\right)\right)+{\mathcal{O}}\left(\sqrt{\frac{d\log(1/\delta)}{\underline{n}}}+\sqrt{\frac{d\log(1/\delta)}{|S|}}\right)$
Theorem B.2 (Main).
Under the assumption of Lemma B.1, we have
$\Delta(S)$ $\leq\text{noise}+\|\bar{\Sigma}_{\text{target}}-\Sigma_{\text{target}}\|\|\Sigma_{S}-\Sigma_{\text{best}}\|_{*}$
+$+\begin{cases}\epsilon_{v*l}^{S}\quad\text{(vision+language)}\\ \left(\epsilon_{v}^{S}+\sqrt{1-\frac{1}{|S|}\sum_{i\in[S]}\langle{\bm{z}}^{v},{\bm{z}}^{l}\rangle)}\right)\quad\text{(vision only)}\end{cases}$
Proof.
Based on Lemma B.1, we will focus on the error cause from selecting subset $S$, that is, $\operatorname{Tr}\Sigma_{\text{target}}\Sigma_{S}$. Since the exact $\Sigma_{\text{target}}$ is unknown, we assume the access to some proxy $\bar{\Sigma}_{\text{target}}$ instead.
Recall that, for any $S$, we have ground-truth $\Sigma_{S}=\mathbb{E}_{{\bm{z}}_{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{l})^{\top}$. Unfortunately, this is not directly observable by the learner. Instead, the learner is able to observe some proxy $\bar{\Sigma}_{S}$ based on the teacher model $\bar{G}_{vl}$ and therefore solving
$\operatorname*{arg\,max}_{S}\operatorname{Tr}\left(\bar{\Sigma}_{\text{target}}\bar{\Sigma}_{S}\right)$
and therefore, denote $\Sigma_{\text{best}}=\operatorname*{arg\,max}_{S^{\prime}\in D_{\text{train}}}\operatorname{Tr}\left(\Sigma_{\text{target}}\Sigma_{S^{\prime}}\right)$
$\operatorname{Tr}\left(\Sigma_{\text{target}}(\Sigma_{\text{best}}-\Sigma_{S})\right)$ $=\operatorname{Tr}\left(\bar{\Sigma}_{\text{target}}(\Sigma_{\text{best}}-\bar{\Sigma}_{S})\right)+\operatorname{Tr}\left(\bar{\Sigma}_{\text{target}}(\bar{\Sigma}_{S}-\Sigma_{S})\right)+\operatorname{Tr}\left((\Sigma_{\text{target}}-\bar{\Sigma}_{\text{target}})(\Sigma_{\text{best}}-\Sigma_{S})\right)$
$\leq\operatorname{Tr}\left(\bar{\Sigma}_{\text{target}}(\bar{\Sigma}_{S}-\Sigma_{S})\right)+\operatorname{Tr}\left((\Sigma_{\text{target}}-\bar{\Sigma}_{\text{target}})(\Sigma_{\text{best}}-\Sigma_{S})\right)$
$\leq\|\Sigma_{\text{target}}\|\|\bar{\Sigma}_{S}-\Sigma_{S}\|_{*}+\|\bar{\Sigma}_{\text{target}}-\Sigma_{\text{target}}\|\|\Sigma_{S}-\Sigma_{\text{best}}\|_{*}$
where the first inequality is by the definition of $\bar{\Sigma}_{S}$ and the second inequality comes from holder’s inequality. Now the key is to upper bound $\|\bar{\Sigma}_{S}-\Sigma_{S}\|_{*}$ based on our chosen strategy.
In option 1, we use the clip embedding from both visual and language modal. That is, choose $\bar{\Sigma}_{S}=\sum_{{\bm{x}}_{vl}\in S}(\bar{G}_{v})^{\top}{\bm{x}}^{v}({\bm{x}}^{l})^{\top}\bar{G}_{l}$. Then we have
$\|\bar{\Sigma}_{S}-\Sigma_{S}\|_{*}\leq\frac{1}{|S|}\|\sum_{{\bm{x}}_{vl}\in S}(\bar{G}_{v})^{\top}{\bm{x}}^{v}({\bm{x}}^{l})^{\top}\bar{G}_{l}-\sum_{{\bm{x}}_{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{l})^{\top}\|_{*}\leq\epsilon_{v*l}^{S}$
In option 2, we use the clip embedding from language model only. That is choose $\bar{\Sigma}_{S}=\sum_{{\bm{x}}_{vl}\in S}\bar{G}_{v}^{\top}{\bm{x}}^{v}({\bm{x}}^{v})^{\top}\bar{G}_{v}$. Then, by definition of $\epsilon_{S}$, we have
$\|\bar{\Sigma}_{S}-\Sigma_{S}\|_{*}$ $\leq\frac{1}{|S|}\|\sum_{{\bm{x}}_{vl}\in S}\bar{G}_{v}^{\top}{\bm{x}}^{v}({\bm{x}}^{v})^{\top}\bar{G}_{v}-\sum_{{\bm{x}}_{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{v})^{\top}\|_{*}+\frac{1}{|S|}\|\sum_{{\bm{x}}_{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{v})^{\top}-\Sigma_{S}\|_{*}$
$\leq\epsilon_{v}^{S}+\frac{1}{|S|}\|\sum_{{\bm{x}}_{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{v})^{\top}-\Sigma_{S}\|_{*}$
Now to further bound the second term, we have
$\frac{1}{|S|}\|\sum_{{\bm{x}}_{vl}\in S}{\bm{z}}^{v}({\bm{z}}^{v})^{\top}-\Sigma_{S}\|_{*}$ $\leq\frac{1}{|S|}\|Z_{v}^{\top}\|_{*}\|Z_{v}-Z_{l}\|_{*}$
$=\frac{1}{|S|}\sqrt{\operatorname{Tr}Z_{v}Z_{v}^{\top}}\sqrt{\operatorname{Tr}(Z_{v}-Z_{l})^{\top}(Z_{v}-Z_{l})}$
$=\frac{1}{|S|}\sqrt{\operatorname{Tr}(I_{n\times n})}\sqrt{2\operatorname{Tr}\left(I_{n\times n}-Z_{v}Z_{l}^{\top}\right)}$
$=\frac{1}{|S|}\sqrt{2|S|(|S|-\sum_{i\in[S]}\langle{\bm{z}}^{v},{\bm{z}}^{l}\rangle)}$
$=\sqrt{1-\frac{1}{|S|}\sum_{i\in[S]}\langle{\bm{z}}^{v},{\bm{z}}^{l}\rangle)}$
Therefore, we finish the proof. ∎
Theorem B.3 (A simplified version of test loss).
Under the assumption that both ${\bm{z}}_{vl},\xi_{vl}$ is zero-mean, maximizing the clip score gap is equivalent to maximize the clip score of the same sample.
${\mathcal{L}}_{\text{target}}(G_{v},G_{l}):=-\mathbb{E}_{{\bm{x}}_{vl}\sim{\mathcal{D}}_{\text{target}}}\langle G_{v}^{\top}{\bm{x}}_{v},G_{l}^{\top}{\bm{x}}_{l}\rangle$
Proof.
For any ${\bm{x}}_{vl}$, we have
$\mathbb{E}_{{\bm{x}}_{vl}^{\prime}\sim{\mathcal{D}}_{\text{target}}}(\langle G_{v}^{\top}{\bm{x}}_{v},G_{l}^{\top}{\bm{x}}_{l}^{\prime}\rangle-\langle G_{v}^{\top}{\bm{x}}_{v},G_{l}^{\top}{\bm{x}}_{l}\rangle)$
$=\langle G_{v}^{\top}{\bm{x}}_{v},G_{l}^{\top}\mathbb{E}_{{\bm{x}}_{vl}^{\prime}\sim{\mathcal{D}}_{\text{target}}}({\bm{x}}_{l}^{\prime}-{\bm{x}}_{l})\rangle$
$=-\langle G_{v}^{\top}{\bm{x}}_{v},G_{l}^{\top}{\bm{x}}_{l}\rangle$
Figure 4: Illustration of different directions for data selection methods for multimodal contrastive learning. Here we use four colors to denote the four main resources we can obtain: CLIP teacher model, downstream target data (which is much smaller than the external multimodal dataset or pretraining dataset), the external image-text dataset, and the external non-CLIP model. Direction 1 denotes the methods that only use the original OAI CLIP teacher model and the downstream target data. Direction 2 represents the methods that use external datasets to train a new CLIP teacher model for improving filtering, like DFN (fang2023data). Direction 3 denotes the methods that use external non-CLIP model to select the data that may be heuristically helpful for downstream tasks, like image without too much text or be more special. In general, D1 method using only CLIP embedding, like negCLIPLoss, is orthogonal to D2. And both D1 and D2 can be combined with D3 to explore better filtering results. In the experiments part of the main paper (Sec. 4), we further show that our proposed D1 methods: NormSim and negCLIPLoss, can outperform all the D3 baselines except the best method “HYPE $\cup$ DFN”. And we can achieve the new state-of-the-art by combining our methods with that method.
Appendix C Illstration of Different Directions for Data Selection for Multimodal Contrastive Learning
We summarize our main idea of categorizing the current top data selection methods in Table 4. In our experiment parts, we consider the baselines below:
D1: OAI CLIP embedding only.
The learner can only access the pretraining dataset (like DataComp-medium), the original OAI CLIP teacher model that is used to extract embeddings, and some target data of the downstream tasks which is much smaller than the pretraining dataset (like ImageNet-1k). In this category, we don’t use any existing external non-CLIP models or any newly trained CLIP model based on external multi-modal dataset. In detail, This category includes (1) CLIPScore (hessel2021clipscore), which only uses CLIPScore for filtering as we mentioned before. (2) Image-based filtering (gadre2023datacomp), which uses ImageNet-1K training data as the downstream target data for data filtering. It applies k-means clustering to the image embeddings of training data and selects clusters closest to the ImageNet-1K embeddings. Gadre et al. (gadre2023datacomp) also try to combine image-based filtering and CLIPScore together. (3) $\mathbb{D}^{2}$ Pruning (maharana2023d2), which represents the dataset as an undirected graph and selects the data by combining difficulty and diversity. They use the CLIP score to initialize their graph.
D2, D3: Accessible external model and multi-modal data.
All the current top baselines enable the learner to utilize external resources, either to train a better CLIP teacher model or to help filtering using existing models’ properies. In detail, (1) DFN (fang2023data) trains another CLIP data filtering network via external high-quality data. Their currently public model (DFN-P) is trained on CC12M (changpinyo2021conceptual12) + CC3M (sharma2018conceptual3) + SS15M (nguyen2022quality), while the best DFN is trained on nonpublic HQITP-350M (fang2023data), which is even larger than DataComp-medium. (2) HYPE (kim2024hype) leverages hyperbolic embeddings (different from CLIP embedding) and the concept of entailment cones to filter out samples with meaningless or underspecified semantics, enhancing the specificity of each sample. (3) HYPE $\cup$ DFN proposed by (kim2024hype) samples subset separately for each method and then merge them. This is the state-of-the-art method on the DataComp benchmark for medium size. (4) Other methods including T-MARS (maini2023t), Devils (yu2023devil), MLM (wang2024finetuned), which leverage external models such as text detection model FAST (chen2021fast), vision-language model (VLM) BLIP-2 (li2023blip) and VLM LLaVA-1.5 (liu2023improved; chiang2023vicuna) to heuristically select the more helpful data, like non-text-only pictures or images with MNIST-style digits. See details in Appendix D.5.
Appendix D Details of Experiments
D.1 Details of Setup
Evaluation.
We measured the model performance on 38 downstream datasets including image classification and retrieval tasks followed by DataComp. The image classification tasks contain ImageNet-1k (deng2009imagenet), ImageNet distribution shifts (wang2019learning; pmlr-v97-recht19a; hendrycks2021natural; hendrycks2021many), 11 datasets from the Visual Task Adaptation Benchmark (VTAB) (zhai2019large) and 3 datasets from WILDS (koh2021wilds; sagawa2021extending). Retrieval datasets contain Flickr30k (young-etal-2014-image), MSCOCO (chen2015microsoft) and WinoGAViL (bitton2022winogavil).
Teacher model architecture.
Our experiments utilize two architectures for OpenAI’s CLIP teacher models: ViT-L/14 and ViT-B/32. Additionally, we use the public version of DFN (DFN-P) proposed by Fang et al. (fang2023data) as a teacher model, and its architecture is also ViT-B/32.
D.2 Computation Cost
Table 4: Comparison of preprocessing time and external resources needed between our method and other D3 category methods. We skip DFN since it’s orthogonal to our negCLIPLoss method and we can directly improve it as mentioned in Table 2. Here since all the baselines below except MLM use a pretrained CLIP model, we only count the time that doesn’t contain that for inferring CLIP image/text embeddings (about 50 L40 hours for OAI CLIP-B/32), which is also adopted in DataComp benchmark (gadre2023datacomp). The external dataset corresponds to the external multimodal dataset used for training or finetuning the external model. Notably, the preprocessing time for the following methods are all approximately linearly proportional to the amount of unfiltered pretrained dataset.
Type Filtering Ext. Model Size of Preprocess Training Avg.
Strategy Used Ext. Dataset Time Time
D1 $\mathbb{D}^{2}$ Pruning NA NA >70 L40 h 65 L40 h 29.5
D3 T-MARS FAST NA 950 L40 h 65 L40 h 34.1
D3 MLM LLaVA-1.5 50k 1120 A100 h 65 L40 h 34.5
D3 Devil fasttext, BLIP-2 NA 510 A100 h 65 L40 h 34.5
D3 HYPE MERU 27M > 120 L40 h 65 L40 h 31.9
D1 Ours1 NA NA 5 L40 h 65 L40 h 35.2
Our algorithm can significantly reduce the computational cost compared to many existing works as shown in Table 4. For example, when the CLIP embeddings are obtained (cost about 50 hours for CLIP-B/32), both T-MARS (maini2023t) and MLM (wang2024finetuned) still require more than 900 hours data pre-processing time to extract the required information from 110M size dataset of DataComp-medium, while we only need about 5 hours. On the other hand, DFN, although has a similar forward speed (i.e. preprocessing time), requires retraining a new CLIP teacher model on the HQITP-350M, which is larger than DataComp-medium.
We give some details in estimating the preprocessing time of other methods:
  • For T-MARS and $\mathbb{D}^{2}$ pruning, we run their official code on DataComp-small (11M) data, and simply scale the preprocessing time by 10 for DataComp-medium, given that the preprocessing time for T-MARS is proportional to the size of the pretraining dataset, while $\mathbb{D}^{2}$ pruning is no faster than linear.
  • For MLM, we get the estimated time from their paper. They mention that they need 6.1 minutes to process 10k samples on A100, which results in 1120 A100 hours for our dataset (110M). We need to mention that their estimation time of calculating CLIP embedding is inaccurate and we can do it much faster than their claim using the DataComp pipeline.
  • For Devil, it needs to run the k-means clustering algorithm from the faiss library on the embedding space, which is estimated to cost 120 L40 hours on DataComp-medium. Using BLIP-2 (li2023blip) to scan the whole dataset will need about 470 A100 hours from the experimental details in (nguyen2023improving). From https://lambdalabs.com/gpu-benchmarks, we roughly assume that 120 L40 hours are at least comparable to 40 A100 hours for K-means clustering.
  • For HYPE, they claim that MERU is as efficient as CLIP, but they still need at least 120 L40 hours for processing 110M data for their final score, since it uses the image embedding clusters on DataComp-medium obtained from running k-means clustering algorithm.
D.3 Details of negCLIPLoss
In negCLIPLoss, we need to get the batch size $|B|$ and the value of the learnable temperature parameter $\tau$ at the final step of the teacher model pretraining stage. For OAI CLIP-L/14 and OAI CLIP-B/32, these values are $\tau=0.01$ and $|B|=32768$. We give the pseudocode of calculating negCLIPLoss in Algorithm 1, which is specially designed for pytorch to accelerate matrix calculation. It can be fully accelerated and the computation cost introduced by the normalization term is negligible compared with the training time or preprocessing time of other top baselines as detailed in Table D.2.
Algorithm 1 negCLIPLoss
  Inputs: image/text embeddings of the pretraining data $F^{vl}=[\{\bar{f}_{vl}(x_{1}^{vl})\},\ldots,\{\bar{f}_{vl}(x_{N}^{vl})\}]^{\top}\in\mathbb{R}^{N\times d}$, batch size $b$, temperature parameter $\tau$, the number of times negCLIPLoss is random $K(=10)$.
  Initialize negCLIPLoss array ${\bm{r}}=[0,\ldots,0]\in\mathbb{R}^{N}$
  for $k=1$ to $K$ do
     Get a random batch division $S_{k}=\{B_{1},\ldots,B_{s}\}$ such that $s=\lceil N/b\rceil$. Every $B_{i}\in S_{k}$ is the index of a batch of data.
     for $j=1$ to $s$ do
        Get batch of embeddings in batch $j$: $F^{vl}_{j}=F^{vl}[B_{j}]\in\mathbb{R}^{b\times d}$
        Get the similarity matrix: $E_{j}=F^{v}_{j}(F^{l}_{j})^{\top}\in\mathbb{R}^{b\times b}$
        Get the CLIPScores: ${\bm{c}}_{j}=\text{diag}(E_{j})\in\mathbb{R}^{b}$
        Define $G_{j}=\exp(E_{j}/\tau)$
        Define ${\bm{g}}^{v}_{j}\in\mathbb{R}^{b}$ be the vector containing the sum of each row vector in $G_{j}$ (i.e., over image).
        Define ${\bm{g}}^{l}_{j}\in\mathbb{R}^{b}$ be the vector containing the sum of each column vector in $G_{j}$ (i.e., over text).
        Get the negCLIPLoss: ${\bm{r}}[B_{j}]={\bm{c}}_{j}-0.5\tau\cdot(\log({\bm{g}}^{v}_{j})+\log({\bm{g}}^{v}_{j}))$, here we use element-wise operation.
     end for
  end for
  Take the mean of each random division as output: $\text{negCLIPLoss}={\bm{r}}/K$
[…]
We use greedy remove methods to achieve the object
$\small S=\arg\max_{|S|=N}\sum_{i\in S}\bar{f}_{v}(x_{i}^{v})^{\top}\left(\frac{1}{|X_{\text{target}}|}\sum_{x_{t}\in X_{\text{target}}}\bar{f}_{v}(x_{t}^{v})\bar{f}_{v}(x_{t}^{v})^{\top}\right)\bar{f}_{v}(x_{i}^{v})$ (15)
when the actual $X_{\text{target}}$ is unknown. In practice, removing one data at a time is too slow. Therefore, we remove a batch of data for every greedy step. In detail, if the number of greedy steps is $\tau$, and let $\bar{\Sigma}_{\text{test},i}=\frac{1}{|S_{i}|}\sum_{j\in S_{i}}\bar{f}_{v}({\bm{x}}_{j}^{v})\bar{f}_{v}({\bm{x}}_{j}^{v})^{\top}$ where $S_{i}$ is the selected subset at step $i$, then we will remove the data satisfies the following equation step-by-step until reaching the final subset size:
$S_{i}\setminus S_{i+1}=\arg\min_{x_{l}\in S_{i}}\left[\bar{f}_{v}(x_{l}^{v})^{T}\cdot\left(\frac{1}{|S_{i}|}\sum_{x_{t}\in S_{i}}\bar{f}_{v}(x_{t}^{v})\bar{f}_{v}(x_{t}^{v})^{\top}\right)\cdot\bar{f}_{v}(x_{l}^{v})\right],\quad i\in\{0,\ldots,\tau-1\}$
[…]
Algorithm 2 NormSim-D with greedy remove strategy
  Inputs: image embeddings of the data after CLIP score filtering $\{\bar{f}_{v}(x_{i}^{v})\}_{i\in S}$, target size $N$, number of greedy steps $\tau$
  Initialize $S_{0}=S,N_{0}=|S|$
  for $t=1$ to $\tau$ do
     Size at step $t$ : $N_{t}=N_{0}-\frac{t}{\tau}(N_{0}-N)$.
     Prior matrix: $\bar{\Sigma}_{\text{test},t-1}=\sum_{j\in S_{t-1}}\bar{f}_{v}(x_{j}^{v})\bar{f}_{v}(x_{j}^{v})^{\top}$
[…]
  end for
D.5 More details in Baselines
We add some details about the baselines used in our paper.
  • Text-based filtering. (gadre2023datacomp) proposes a text-based filtering that tries to select the data that contains caption overlapping with the class name from ImageNet-21K or ImageNet-1K.
  • Image-based filtering. (gadre2023datacomp) also proposes a heuristic way to sample the visual content overlaps with ImageNet-1K classes. They first apply filtering by language (only choose English caption by fasttext (joulin2016bag)) and caption length (over two words and 5 characters). Then they cluster the image embeddings from training data to 100K groups using Faiss (johnson2019billion), and keep the groups whose cluster center is the nearest neighbor to at least one image embedding of ImageNet-1K image.
  • $\mathbb{D}^{2}$ Pruning. (maharana2023d2) tries to represent the dataset as an undirected graph for coreset selection. They assign the difficulty for each example and use message passing to update the difficulty score incorporating the difficulty of its neighboring examples, and finally try to keep both diverse and difficult subsets. For our experiments, we adhere to the default hyperparameters of $\mathbb{D}^{2}$ on DataComp as specified in their official codebase.
  • T-MARS (maini2023t) uses a text detection model like FAST (chen2021fast) to filter out the data that only contain the texts of caption in the image and don’t have other useful image features.
  • Devils (yu2023devil) combines many ways for data filtering. At the very first it filter data based on heuristic rules like text length, frequency of texts, and image size, and it also use CLIPScore for cross-modality matchment. Then it adopts target distribution alignment methods similar to image-based filtering, but instead of using ImageNet-1k only, it uses 22 downstream tasks as the target set. Further, it adopts external models fasttext (joulin2016bag) to remove non-English captions and image-captioning model BLIP-2 (nguyen2024improving) to select images with MNIST-style digits.
  • MLM (wang2024finetuned) prompts GPT-4V to construct instruction data including the image-text data, and use it to fine-tune a smaller vision-language model like LLaVA-1.5 (liu2023improved; chiang2023vicuna) into a filtering network. Nevertheless, the number of parameters of LLaVA-1.5 is still much larger than CLIP, and thus LLaVA-1.5 has a much longer preprocessing time as mentioned in Table D.2.
D.6 Hyperparameters
The main hyper-parameters of our negCLIPLoss and NormSim are the target numbers for filtering, which is also the main concerns for all the top baselines like DFN, MLM, and T-MARS. However, from the current paper about the scaling law for data filtering (goyal2024scaling), the downsampling size should depend on the computation budget. When you have more budget, you should sample more data for learning. And we can use their fitting formula to get some good downsampling ratios.
In the case of the DataComp benchmark, however, we can set the sampling ratio as some previous baselines. but in general, for NormSim, we can also filter out the data until some train-dataset-independent threshold since this score only depends on the norm $p$ and target data. For example, we find it good to set the threshold as 0.7 for $\text{NormSim}_{\infty}$(Target) and 0.15 for $\text{NormSim}_{2}$(IN-1k). For negCLIPLoss, we can find a similar way: first find the ratio of data whose CLIP score is higher than 0.21, then use this ratio to filter negCLIPLoss.
Besides, we found an effective parameter tuning way particularly for data selection problems, that is, visualization. First randomly selecting a small subset (like 1000 data) on some pretraining data subset and calculating the target metric on them, then visualizing the data with different scores, it turns out to be a very effective way to give some guidance on how to select the downsampling ratios.
Appendix E Additional Results
E.1 Stability Analysis of Batch Sampling Numbers in negCLIPLoss
We show that negCLIPLoss is not sensitive to the number of random select batches $K$ in Figure 5.
Figure 5: Results of negCLIPLoss with a different number of batch samples (denoted as $K$) on DataComp-medium. Solid lines denote negCLIPLoss, while dashed lines denote CLIPScore. Here, we use OAI CLIP-L/14 as the pretrained model. We can see that once $K\geq 5$, negCLIPLoss consistently outperforms CLIPScore across all subtask metrics. In the main paper, we set $K=10$.
E.2 Complete Comparison Results for D1 categories
[…]
Table 5: Results on DataComp-medium from methods that use only OpenAI’s CLIP-L/14 model, i.e., all methods are from the D1 category. The “dataset size” represents the size of the subset obtained from different approaches. NormSim(IN-1k) denotes using the training data of ImageNet-1k as the target while NormSim(Target) represents using that of all 24 available downstream tasks. NormSim-D refers to the methods that use an iteratively selected subset from the training set as the target proxy.
Filtering Strategy Dataset IN-1k IN Shift VTAB Retr. Avg.
Size (1 task) (5) (11) (3) (38)
No filtering (gadre2023datacomp) 110M 17.3 15.0 25.2 21.3 25.6
CLIPScore (20%) (hessel2021clipscore) 22M 25.4 22.7 31.8 22.0 31.0
CLIPScore (30%) (hessel2021clipscore) 33M 26.4 23.6 32.6 24.5 32.2
Image-based (gadre2023datacomp) 24M 25.5 21.9 30.4 24.6 29.9
CLIPScore (30%) $\cap$ Image-based 11M 27.4 23.9 31.9 21.4 30.8
$\mathbb{D}^{2}$ Pruning (maharana2023d2) 22M 23.2 20.4 31.4 18.7 29.5
negCLIPLoss (20%) 22M 27.4 23.8 33.7 23.7 32.5
negCLIPLoss (30%) 33M 27.9 24.6 33.2 25.1 32.9
[…]
CLIPScore (30%) $\cap$ $\text{NormSim}_{2}$(IN-1k) 22M 29.1 25.4 35.8 24.1 33.4
CLIPScore (30%) $\cap$ $\text{NormSim}_{2}$(Target) 22M 28.9 25.1 32.7 23.6 32.5
CLIPScore (30%) $\cap$ $\text{NormSim}_{\infty}$(IN-1k) 22M 29.7 25.9 33.7 24.1 33.7
CLIPScore (30%) $\cap$ $\text{NormSim}_{\infty}$(Target) 22M 30.2 26.2 35.0 23.4 33.9
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{2}$(IN-1k) 22M 30.4 26.4 35.4 25.6 34.3
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{2}$(Target) 22M 30.6 26.2 35.2 25.5 33.9
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(IN-1k) 22M 31.9 27.3 34.8 25.0 34.4
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target) 22M 31.7 27.2 36.0 26.0 35.0
E.3 Universality of negCLIPLoss over Different Teacher Models
We show the complete results of applying our methods to different teacher models like OAI CLIP-B/32 and DFN-P in Table 6. Detail descriptions are in Sec. 4.
Table 6: Results on DataComp-medium from the top methods that use only OpenAI’s CLIP-B/32 model or public version of DFN (DFN-P).
OAI CLIP-B/32 Dataset IN-1k IN Shift VTAB Retr. Avg.
Size (1) (5) (11) (3) (38)
CLIPScore (20%) 22M 27.0 23.8 33.0 22.9 32.2
CLIPScore (30%) 33M 27.6 24.2 33.6 25.1 33.2
negCLIPLoss (20%) 22M 28.9 24.8 34.3 24.3 33.0
negCLIPLoss (30%) 33M 28.8 25.1 33.7 26.6 33.6
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target) 22M 32.4 27.4 35.9 26.3 35.2
DFN-P
CLIPScore (15%) 16M 25.9 23.3 32.9 21.9 31.6
CLIPScore (17.5%) 19M 30.2 26.8 34.1 26.5 33.8
CLIPScore (20%) 22M 29.7 26.8 33.0 27.0 33.1
CLIPScore (30%) 33M 28.4 24.7 33.2 26.8 32.7
negCLIPLoss (15%) 16M 31.3 27.3 35.8 26.4 34.6
negCLIPLoss (17.5%) 19M 31.2 27.5 35.7 27.0 34.7
negCLIPLoss (20%) 22M 30.7 27.4 33.6 27.5 33.8
negCLIPLoss (30%) 33M 28.9 25.5 33.4 27.3 33.2
negCLIPLoss (30%) $\cap$ $\text{NormSim}_{\infty}$(Target) 22M 29.4 23.6 33.5 24.2 32.5
negCLIPLoss (17.5%) $\cap$ $\text{NormSim}_{\infty}^{\text{B/32}}$(Target) 16M 31.6 27.3 37.2 25.5 35.7
E.4 $\text{NormSim}_{\infty}$ is Better than Nearest Neighbor Selection
We also try to use near-neighbor selection for aligning downstream distribution. Here, we calculate the ranks of pretraining data for each target (the higher the rank, the higher the similarity), and then for each pre-train data, we keep its highest rank. Finally, we select the data with the highest ranks as the nearest neighbor selected subset.
In Table 7, we show that given the training data of 22 downstream tasks, our $\text{NormSim}_{\infty}$ can outperform near neighbor selection under the same downsampling ratio. The reason may be that the distribution between the target and pretraining set is not well aligned, so if you force the algorithm to find the nearest train data for each target, that train data may be sometimes random and not helpful. On the other hand, $\text{NormSim}_{\infty}$ will not select this kind of data. It will select the data whose best similarity score exceeds some general threshold, rather than just consider ranks.
Table 7: Comparison between $\text{NormSim}_{\infty}$ and nearest neighbor selection. We use OAI CLIP-L/14 as the teacher model and assume both methods have been intersected with negCLIPLoss (30%). The size of the selected subset is 22M.
Filtering Strategy IN-1k VTAB Avg.
negCLIPLoss (30%) 27.9 33.2 32.9
Nearest Neibor Selection 31.5 34.9 34.0
$\text{NormSim}_{\infty}$(Target) 31.7 36.0 35.0
E.5 Vision-Only NormSim is Better than Using Both Vision and Language
[…]
Table 8: Ablation Study on the NormSim and its variants on DataComp-small (11M). All experiments first select 45% data based on the CLIP score, then use corresponding approaches to obtain 3.3M data.“image” or “text” means using the variance of image or text embeddings to represent $\bar{\Sigma}_{\text{target}}$, and “image $\times$ text” means representing $\bar{\Sigma}_{\text{target}}$ with the cross-covariance of image and text embeddings.
Filtering Strategy $\cap$ CLIP score (45%) IN-1k IN Dist. Shift VTAB Retrieval Average
Random Sampling 4.2 4.9 17.2 11.6 15.6
NormSim (IN-1k, image) 5.2 5.5 19.0 12.2 17.4
NormSim (IN-1k, text) 3.9 4.2 16.3 11.3 14.9
NormSim (IN-1k, image $\times$ text) 4.3 4.9 17.5 11.8 15.9
NormSim-D (image) 4.7 5.4 19.7 11.7 17.3
NormSim-D (text) 3.5 4.1 16.7 11.1 15.4
NormSim-D (image $\times$ text) 3.6 4.2 18.4 11.1 15.8
Appendix F Additional Visualization
We further visualize more data with different negCLIPLoss in Figure 6, 7 and 8. And similar for $\text{NormSim}_{\infty}$(Target) in Figure 9, 10 and 11.
Figure 6: Visualization of a small subset whose negCLIPLoss rank top 100% high in DataComp-medium.
Figure 7: Visualization of a small subset whose negCLIPLoss rank top 50% high in DataComp-medium.
Figure 8: Visualization of a small subset whose negCLIPLoss rank top 10% high in DataComp-medium.
Figure 9: Visualization of the images from a small subset whose $\text{NormSim}_{\infty}$(Target) rank top 100% high in DataComp-medium.
Figure 10: Visualization of the images from a small subset whose $\text{NormSim}_{\infty}$(Target) rank top 50% high in DataComp-medium.
Figure 11: Visualization of the images from a small subset whose $\text{NormSim}_{\infty}$(Target) rank top 10% high in DataComp-medium.